# library.bib

### v1

@@ I Introduction
Reinforcement learning (RL) policies showed promise in addressing these challenges, partly because they are trained in mature high-fidelity simulators with whole-body dynamics and kinematics. However, RL requires vast amounts of data.
@@ I-B Solution Methods for Real-Time NMPC of Legged Robots
Roboticists have developed various strategies to address the computational burden of whole-body NMPC. The most common solution methods for NMPC include sequential quadratic programming (SQP) [nocedalNumericalOptimization2006] and differential dynamic programming (DDP) [mayneSecondorderGradientMethod1966], both of which solve a local convex subproblem at each iteration. Solving NMPC to optimality can be prohibitively slow; therefore, only one or a few iterations are typically performed online at each control instant [diehlRealTimeIterationScheme2005]. Furthermore, specialized matrix factorization methods that exploit the temporal structure of the problem (e.g., Riccati recursion), specialized algorithms for analytical derivatives [carpentierAnalyticalDerivativesRigid2018, katayamaEfficientSolutionMethod2021, katayamaModelPredictiveControl2023] and parallelization [jalletParallelProximalLinearQuadratic2024, katayamaEfficientSolutionMethod2021] enabled real-time NMPC on hardware for quadrupeds [neunertWholeBodyNonlinearModel2018, sleimanUnifiedMPCFramework2021, grandiaPerceptiveLocomotionNonlinear2023, mastalliInverseDynamicsMPCNullspace2023] and humanoids [koenemannWholebodyModelpredictiveControl2015, dantecWholeBodyModelPredictive2022].
@@
Another augmented Lagrangian technique for solving QP subproblems is the alternating direction method of multipliers (ADMM) [boydDistributedOptimizationStatistical2010]. ADMM is a general framework that splits optimization problems into smaller subproblems that are easier to solve and has gained popularity in robotics [budhirajaDynamicsConsensusCentroidal2019, meduriBiConMPNonlinearModel2023, bravo-palaciosLargeScaleADMMbasedCoDesign2022, aydinogluRealTimeMultiContactModel]. In particular, the operator splitting QP solver OSQP leverages ADMM to handle linear equality and inequality constraints while only requiring a single sparse matrix factorization followed by multiple cheap iterations [stellatoOSQPOperatorSplitting2020]. Yet, there is currently no demonstration of its applicability for solving the QP subproblems arising in whole-body NMPC of legged robots in real time, even though some SQP implementations using OSQP exist [anderssonCasADiSoftwareFramework2019, verschuerenAcadosModularOpensource2022] and have been deployed on lower-dimensional robot manipulators [jordanaStagewiseImplementationsSequential]. While ADMM may need many iterations to provide high-accuracy solutions, it can rapidly converge to low-accuracy solutions and be warm-started, making it a practical option for solving very complex NMPC problems in real time.
@@ I-C Contributions
The main contributions of this paper are threefold. First, we present a whole-body NMPC formulation that leverages ADMM to solve the inner QP subproblem of an SQP method with inequality constraints in real time. Through extensive simulation, we demonstrate the advantages of rapid, low-accuracy solutions obtained from ADMM by benchmarking the impact of solution accuracy on closed-loop performance in the presence of unavoidable inertial modeling errors, dynamics discretization errors, and computation delay.
@@ II-A Background on Non-linear MPC
where ${\bm{\mathcal{X}}}$ is the set of decision variables composed of the states $\bm{x}_{k}$ and control inputs $\bm{u}_{k}$ at stage $k\in\{0,...N\}$, $\hat{\bm{x}}$ is the measured state, $\bm{\theta}$ is a parameter vector, $l(\cdot)$ is the running cost, $l_{N}(\cdot)$ is the terminal cost, $\bm{f}(\cdot)$ is the distrete-time dynamics, $\bm{g}(\cdot)$ and $\bm{g}_{N}(\cdot)$ are the path equality and inequality constraints.
@@ II-B1 Decision Variables
$\bm{\mathcal{X}}=\{\{\bm{q}_{k},\bm{v}_{k}\}_{k=0}^{N},\{\bm{\tau}_{k},\bm{F}_{c,k}\}_{k=0}^{N-1}\}.$ (2)
@@ II-B2 Cost Function
We use a simple quadratic cost function that penalizes the deviation from constant velocity and torso height commands along the NMPC horizon. We use simple regularization for all other decision variables. The difference between the decision variable $\bm{\tau}_{0}$ and its value from the previous solution is penalized to eliminate vibrations on hardware. The cost function has a least-squares form. Therefore, we approximate the Hessian of the Lagrangian with the Gauss-Newton Hessian to reduce solve time [diehlRealTimeIterationScheme2005].
@@ II-B3 Dynamics Constraints
where $\bm{J}_{c}\in\mathbb{R}^{3\times(n_{j}+6)}$ is the Jacobian of contact point $c$. The wrench on the floating base is zeroed, and $\bm{\tau}_{k}$ corresponds to the last $n_{j}$ rows in (4). The acceleration $\dot{\bm{v}}_{k}$ is computed from finite differences:
@@ II-B4 Contact and Swing Constraints
The NMPC takes as input a contact sequence $\theta_{c,k}\in\left\{0,1\right\}$ for each contact point $c$ and stage $k$, where $\theta_{c,k}=1$ in contact and $\theta_{c,k}=0$ in swing. Contact forces are constrained to $\bm{0}_{3\times 1}$ in swing and to the friction cone in stance:
$(1-\theta_{c,k})\bm{F}_{c,k}=\bm{0}_{3\times 1}$ (6)
$\theta_{c,k}\mu F_{c,z,k}\geq\theta_{c,k}\sqrt{F_{c,x,k}^{2}+F_{c,y,k}^{2}},$ (7)
where $F_{c,x,k}$ and $F_{c,y,k}$ are the world-frame tangential forces in the forward and lateral directions. The multiplication by $(1-\theta_{c,k}\phantom{})$ and $\theta_{c,k}$ respectively deactivates the constraints in stance and swing without removing them from the problem at runtime. This keeps the sparsity structure of the QP subproblems fixed, which facilitates warmstart and avoids repeating expensive symbolic factorizations.
@@
$\theta_{c,k}\bm{J}_{c,xy,k}\dot{\bm{q}}_{k}=\bm{0}_{2\times 1},$ (9)
where $\bm{J}_{c,xy,k}$ maps $\dot{\bm{q}}_{k}$ to the velocity tangent to the ground. We use the adaptive segmentation algorithm detailed in [bledtImplementingRegularizedPredictive2019] to segment each $\theta_{c,k}$ and $dt_{k}$ to match the specified gait.
@@ II-C Solution Method
The NMPC formulation is expressed in the form of the general non-linear program (12)
@@
where $\beta_{i}$ is obtained from the linesearch algorithm from [grandiaPerceptiveLocomotionNonlinear2023].
Solving (LABEL:eq:sqp_step) is the most computationally expensive step, especially when inequality constraints are present. In this work, we solve (LABEL:eq:sqp_step) using the ADMM algorithm from the QP solver OSQP [stellatoOSQPOperatorSplitting2020], which provides low-accuracy solutions within a few cheap iterations and can easily be wamstarted. This allows to tailor the solution accuracy to the computational budget by running a small but fixed number of iterations at high rates. The feasibility of the self-collision avoidance constraint (10) is enhanced by applying the CBF constraint (11) at the first timestep. Overall, our approach allows to enforce equality and inequality constraints in real time with sufficient accuracy for hardware deployment.
@@ II-D2 NMPC Implementation for the MIT Humanoid
We run a single SQP iteration at every control instant, and the QP solver runs a fixed number of iterations. This allows to update solutions at deterministic rates and facilitates prediction delay compensation, which is performed by linearly interpolating the solved trajectory forward in time.
@@ III Simulation Experiments
The goal of the simulation experiments is to assess the effect of the solution accuracy of the QP subproblem (LABEL:eq:sqp_step) on closed-loop performance considering three unavoidable sources of error: inertial modeling errors, dynamics discretization errors, and computation delay. Additionally, we evaluate the impact of CBFs on self-collision avoidance.
@@ III-A Experimental Setup
The NMPC controller employs an approximate contact model with coarse dynamics integration timesteps as detailed in Section II-B, but is evaluated in a custom high-fidelity simulator. This simulator strictly enforces contact complementarity constraints, friction cone, and actuator torque-speed limits with a fine simulation timestep of 0.50.50.5~{}0.5ms.
To thoroughly evaluate a controller’s closed-loop performance, we conduct large-scale simulations comprising 500 individual runs, each lasting 6 seconds. During each simulation, the robot is initialized in the same standing pose. It starts walking while subjected to a constant velocity command $\bm{v}_{\mathrm{cmd}}=[v_{x,\mathrm{cmd}},v_{y,\mathrm{cmd}},\omega_{z,\mathrm{cmd}}]\in[-3,3]~{}\mathrm{m/s}\times[0,2]~{}\mathrm{m/s}\times[0,5]~{}\mathrm{rad/s}$, where $v_{x,\mathrm{cmd}}$, $v_{y,\mathrm{cmd}}$, and $\omega_{z,\mathrm{cmd}}$ are the forward, lateral, and turning velocity commands, respectively. A simulation is deemed successful if the robot avoids falling and satisfies the simple self-collision constraints (Fig. 2a) throughout the entire duration.
@@
For each controller tested, we analyse the number of successful simulations, normalized by the maximum number of successes obtained by the most successful controller (i.e. with $dt=20$ ms at 150 ADMM iterations and without inertial modeling error.
@@ III-B Effect of Inertial Modeling Errors
Fig. 3a shows the number of successful simulations against the number of ADMM iterations for increasing inertial modeling errors. The number of successes increases with the number of ADMM iterations. However, inertial modeling errors render higher-accuracy solutions less effective. Fig. 3b shows the difference between the number of successes at 20 ADMM iterations and the maximum number of successes from Fig. 3a for each level of inertial model randomization. This assesses the potential success improvement obtainable by increasing solution accuracy when computation delay is neglected. The potential for success improvement reduces as modeling errors increase, indicating that the real MIT Humanoid may not benefit significantly from higher solution accuracy. These effects are expected to be more pronounced on hardware. For example, our simulator does not account for foot rolling contact, backlash, transmission compliance, torque-dependent friction [wangDirectionalEfficiencyGeared2015], sensor noise, or low-level current control.
Figure 3: a) Normalized number of successes for increasing ADMM iterations and inertial modeling errors without CBFs and computation delay. b) Potential improvement between 20 ADMM iterations and maximum success as a function of the inertia randomization level of the simulated model.
@@ III-C Effect of Dynamics Discretization Error
We draw similar observations as in Section III-B for the effects of dynamics discretization errors. Fig. 4a shows the normalized number of successful simulations against the number of ADMM iterations. Between 20 to 40 iterations, the controllers with $dt\in\{20,25,30\}$ ms perform identically, and it takes many more iterations for $dt=20$ ms to outperform the coarser timesteps.
The formulations with larger $dt$ benefit less from increasing solution accuracy. As in Fig. 3b, Fig. 4b shows the potential success improvement for each discretization timestep. The potential success improvement decreases for larger timesteps, indicating that models with larger discretization errors, typically used for real-time NMPC, may not benefit as much from increased solution accuracy when used in a high-fidelity simulator and on hardware. For instance, the least accurate controller that is still stable on hardware ($dt=50$ ms) does not benefit from running more than 20 ADMM iterations. Even though it has the lowest number of successes, it has still been successfully deployed on hardware for a walking speed of up to 0.8 m/s.
Figure 4: a) Normalized number of successes for increasing ADMM iterations and integration timestep without CBFs and computation delay. b) Potential improvement between 20 ADMM iterations and maximum success as a function of the discretization timestep for a fixed time horizon of 0.8 s.

### v2

@@ Abstract
Index Terms:
Legged Robots, Whole-Body Motion Planning and Control, Optimization and Optimal Control, Solution Accuracy, Self-Collision Avoidance
@@ I Introduction
Reinforcement learning (RL) policies showed promise in addressing these challenges, partly because they are trained in mature high-fidelity simulators with whole-body dynamics and kinematics. However, RL requires vast amounts of data and cannot easily enforce system constraints.
@@ I-B Solution Methods for Real-Time NMPC of Legged Robots
Various strategies are used to address the computational burden of whole-body NMPC. Common solution methods include sequential quadratic programming (SQP) [nocedalNumericalOptimization2006] and differential dynamic programming (DDP) [mayneSecondorderGradientMethod1966], both of which solve a local convex subproblem at each iteration. Solving NMPC to optimality can be prohibitively slow; therefore, only one or a few iterations are typically performed online at each control instant [diehlRealTimeIterationScheme2005]. Furthermore, specialized matrix factorization methods that exploit the temporal structure of the problem (e.g., Riccati recursion), specialized algorithms for analytical derivatives [carpentierAnalyticalDerivativesRigid2018, katayamaEfficientSolutionMethod2021, katayamaModelPredictiveControl2023] and parallelization [jalletParallelProximalLinearQuadratic2024, katayamaEfficientSolutionMethod2021] enabled real-time NMPC on hardware for quadrupeds [neunertWholeBodyNonlinearModel2018, sleimanUnifiedMPCFramework2021, grandiaPerceptiveLocomotionNonlinear2023, mastalliInverseDynamicsMPCNullspace2023] and humanoids [koenemannWholebodyModelpredictiveControl2015, dantecWholeBodyModelPredictive2022].
@@
Another augmented Lagrangian technique for solving QP subproblems is the alternating direction method of multipliers (ADMM) [boydDistributedOptimizationStatistical2010]. ADMM is a general framework that splits optimization problems into smaller subproblems that are easier to solve and has gained popularity in robotics [budhirajaDynamicsConsensusCentroidal2019, meduriBiConMPNonlinearModel2023, bravo-palaciosLargeScaleADMMbasedCoDesign2022, aydinogluRealTimeMultiContactModel2022]. In particular, the operator splitting QP solver OSQP leverages ADMM to handle linear equality and inequality constraints while only requiring a single sparse matrix factorization followed by multiple cheap iterations [stellatoOSQPOperatorSplitting2020]. Yet, there is currently no demonstration of its applicability for solving the QP subproblems arising in whole-body NMPC of legged robots in real time, even though some SQP implementations using OSQP exist [anderssonCasADiSoftwareFramework2019, verschuerenAcadosModularOpensource2022] and have been deployed on lower-dimensional robot manipulators [jordanaStagewiseImplementationsSequential2023]. While ADMM may need many iterations to provide high-accuracy solutions, it can rapidly converge to low-accuracy solutions and be warm-started, making it a practical option for solving very complex NMPC problems in real time.
In this paper, we present a whole-body NMPC formulation and a solver that leverages ADMM to solve the inner QP subproblem of an SQP method with inequality constraints in real time.
@@ I-C Contributions
The main contributions of this paper are threefold. First, we demonstrate the advantages of rapid, low-accuracy QP subproblem solutions obtained from ADMM by benchmarking the impact of solution accuracy on closed-loop performance in the presence of unavoidable inertial modeling errors, dynamics discretization errors, and computation delay.
@@ II-A Background on Non-linear MPC
where ${\bm{\mathcal{X}}}$ is the set of decision variables composed of the states $\bm{x}_{k}$ and control inputs $\bm{u}_{k}$ at stage $k\in\{0,...N\}$, $\hat{\bm{x}}$ is the measured state, $\bm{\theta}$ is a parameter vector, $l(\cdot)$ is the running cost, $l_{N}(\cdot)$ is the terminal cost, $\bm{f}(\cdot)$ is the discrete-time dynamics, $\bm{g}(\cdot)$ and $\bm{g}_{N}(\cdot)$ are the path equality and inequality constraints.
@@ II-B1 Decision Variables
$\bm{\mathcal{X}}:=\left\{\left\{\bm{q}_{k},\bm{v}_{k}\right\}_{k=0}^{N},\left\{\bm{\tau}_{k},\left\{\bm{F}_{c,k}\right\}_{\forall c\in\mathcal{C}}\right\}_{k=0}^{N-1}\right\}.$ (2)
@@ II-B2 Cost Function
The running cost is given by
$l=\left\lVert\delta\bm{v}_{k}\right\rVert_{\bm{W}_{\bm{v}}}^{2}+\left\lVert\delta\bm{q}_{k}\right\rVert_{\bm{W}_{\bm{q}}}^{2}+\textstyle\sum_{c\in\mathcal{C}}\left\lVert\delta\bm{F}_{c,k}\right\rVert_{\bm{W}_{\bm{F}_{c}}}^{2},$ (3)
[…]
@@ II-B3 Dynamics Constraints
where $\bm{J}_{c}\in\mathbb{R}^{3\times(n_{j}+6)}$ is the Jacobian of contact point $c$. The RNEA algorithm and its derivatives are computationally cheaper to compute than the forward dynamics, and can advantageously increase the sparsity of the NMPC [katayamaEfficientSolutionMethod2021]. The wrench on the floating base is zeroed, and $\bm{\tau}_{k}$ corresponds to the last $n_{j}$ rows in (5). The acceleration $\dot{\bm{v}}_{k}$ is computed from finite differences:
@@ II-B4 Contact and Swing Constraints
The NMPC takes as input a contact sequence $\gamma_{c,k}\in\left\{0,1\right\}$ for each contact point $c$ and stage $k$, where $\gamma_{c,k}=1$ in contact and $\gamma_{c,k}=0$ in swing. Contact forces are constrained to $\bm{0}_{3\times 1}$ in swing and to the friction cone in stance:
$(1-\gamma_{c,k})\bm{F}_{c,k}=\bm{0}_{3\times 1}$ (7)
$\gamma_{c,k}\mu F_{c,z,k}\geq\gamma_{c,k}\sqrt{F_{c,x,k}^{2}+F_{c,y,k}^{2}},$ (8)
where $F_{c,x,k}$ and $F_{c,y,k}$ are the world-frame tangential forces in the forward and lateral directions. The multiplication by $(1-\gamma_{c,k}\phantom{})$ and $\gamma_{c,k}$ respectively deactivates the constraints in stance and swing without removing them from the problem at runtime. This keeps the sparsity structure of the QP subproblems fixed, which facilitates warmstart and avoids repeating expensive symbolic factorizations.
@@
$\gamma_{c,k}\bm{J}_{c,xy,k}\dot{\bm{q}}_{k}=\bm{0}_{2\times 1},$ (10)
where $\bm{J}_{c,xy,k}$ maps $\dot{\bm{q}}_{k}$ to the velocity tangent to the ground. We use the adaptive segmentation algorithm detailed in [bledtImplementingRegularizedPredictive2019] to segment each $\gamma_{c,k}$ and $dt_{k}$ such that the predicted gait events (transitions from stance to swing and from swing to stance) match the predicted gait along the prediction horizon.
@@ II-C Solution Method
The NMPC formulation is solved with a custom general-purpose solver that exploits sparsity. Alternatively, a solver tailored for optimal control, as in [jordanaStagewiseImplementationsSequential2023] could be used by defining $\bm{x}_{k}=[\bm{q}_{k},\bm{v}_{k}]^{\top}$ and $\bm{u}_{k}=[\bm{\tau}_{k},\{\bm{F}_{c,k}\}_{\forall c\in\mathcal{C}},\bm{v}_{k+1}]^{\top}$, and recasting the problem into the form of (1).
The NMPC is expressed in the form of the general non-linear program (13)
@@
where $\beta_{i}$ is obtained from a linesearch algorithm.
Solving (LABEL:eq:sqp_step) is the most computationally expensive step, especially when inequality constraints are present. In this work, we solve (LABEL:eq:sqp_step) using the ADMM algorithm from the QP solver OSQP [stellatoOSQPOperatorSplitting2020], which provides low-accuracy solutions within a few cheap iterations and can easily be wamstarted. This allows to tailor the solution accuracy to the computational budget by running a small but fixed number of iterations at high rates.
@@ II-D2 NMPC Implementation for the MIT Humanoid
We run a single SQP iteration at every control instant, utilizing the previous unshifted solution as an initial guess. The inner OSQP solver runs a fixed number of iterations. This allows to update solutions at deterministic rates and facilitates prediction delay compensation, which is performed by linearly interpolating the solved trajectory forward in time. The default sparse QDLDL linear solver is used to efficiently exploit the problem’s temporal sparsity and the robot’s kinematics and dynamics sparsity. To reduce hardware vibrations and avoid additional time-consuming matrix factorizations, the adaptive_rho option in OSQP is disabled. The linesearch algorithm from [grandiaPerceptiveLocomotionNonlinear2023] is used as it also helped reduce vibrations on hardware.
@@ II-D3 Low-level Control
II-D4 State Estimation
A linear Kalman filter fuses contact state estimates, joint encoders and accelerometer measurements to estimate the linear velocity of the torso, as in [bledtMITCheetahDesign2018]. The contact states of the heels and toes are assumed to match the pre-specified gait. This strategy was sufficient to walk stably on uneven terrain as shown in Section IV.
@@ III Simulation Experiments
The simulations aim to assess the effect of the solution accuracy of the QP subproblem (LABEL:eq:sqp_step) on closed-loop performance considering three sources of error: inertial errors, dynamics discretization errors, and computation delay. We also evaluate the impact of CBFs on self-collision avoidance.
@@ III-A Experimental Setup
The NMPC uses an approximate contact model with coarse dynamics integration timesteps, but is evaluated in a custom high-fidelity simulator [CheetahSoftware2024]. This simulator enforces contact complementarity constraints, friction cone, and actuator torque-speed limits with a fine timestep of 0.50.50.5~{}0.5ms.
To evaluate a controller’s closed-loop performance, we conduct large-scale simulations comprising 500 individual runs, each lasting 6 seconds. During each simulation, the robot is initialized in the same standing pose. It starts walking while subjected to a constant velocity command $\bm{v}_{\mathrm{cmd}}=[v_{x,\mathrm{cmd}},v_{y,\mathrm{cmd}},\omega_{z,\mathrm{cmd}}]\in[-3,3]~{}\mathrm{m/s}\times[0,2]~{}\mathrm{m/s}\times[0,5]~{}\mathrm{rad/s}$, where $v_{x,\mathrm{cmd}}$, $v_{y,\mathrm{cmd}}$, and $\omega_{z,\mathrm{cmd}}$ are the forward, lateral, and turning velocity commands, respectively. A simulation is deemed successful if the robot avoids falling and satisfies the simple self-collision constraints (Fig. 2a) throughout the entire duration. Every controller experiences the same set of 500 velocity commands selected randomly.
@@
For each controller tested, we analyse the number of successful simulations, normalized by the maximum number of successes obtained by the most successful controller (i.e. 255 successes for $dt=20$ ms at 150 ADMM iterations, without inertial modeling error).
@@ III-B Effect of Inertial Modeling Errors
Fig. 3 shows the number of successful simulations against the number of ADMM iterations for increasing inertial modeling errors, for $dt=25$ ms. The number of successes increases with the number of ADMM iterations.
However, inertial modeling errors render higher-accuracy solutions less effective. We computed the difference between the number of successes at 20 ADMM iterations and the maximum number of successes for each level of inertial model randomization in Fig. 3. This assesses the potential success improvement obtainable by increasing solution accuracy when computation delay is neglected. The potential for success improvement respectively reduces from 51% to 46, 32, 27 and 21% for inertial errors of $\pm$ 0, 10, 20, 30 and 40%, indicating that the real MIT Humanoid may not benefit significantly from higher solution accuracy. These effects should be more pronounced on hardware, since our simulator does not model foot rolling contact, backlash, transmission compliance, sensor noise, or low-level current control.
Figure 3: Normalized number of successes for increasing ADMM iterations and inertial modeling errors without CBFs and computation delay for $dt$ = 25 ms. The potential success improvement decreases as the inertia randomization increases (i.e., 51, 46, 32, 27 and 21%, respectively).
@@ III-C Effect of Dynamics Discretization Error
Fig. 4 shows the normalized number of successful simulations against the number of ADMM iterations for increasing discretization timesteps. Between 20 to 40 iterations, the controllers with $dt\in\{20,25,30\}$ ms perform identically, and it takes many more iterations for $dt=20$ ms to outperform the coarser timesteps.
The potential success improvement respectively decreases from 67% to 51, 41, 18, 7 and 1% as timesteps increase from 20 ms to 25, 30, 35, 40 and 50 ms, indicating that models with larger discretization errors, typically used for real-time NMPC, may not benefit as much from increased solution accuracy when used in a high-fidelity simulator and on hardware. For instance, the controller with $dt=50$ ms does not benefit from running more than 20 ADMM iterations. Despite being the least successful, it made the real robot walk at up to 0.8 m/s.
Figure 4: Normalized number of successes for increasing ADMM iterations and integration timestep without CBFs and computation delay. The potential success improvement decreases as the integration timestep increases (i.e., 67, 51, 41, 18, 7 and 1%, respectively).
@@ III-D Effect of Computation Delay
Surprisingly, the number of successes is 17% higher with delay for 30 ADMM iterations. We found that the low-level PD control law (16) becomes more effective under delay when paired with the prediction delay compensation strategy. Without delay, the feedback terms have minimal impact, because the setpoints from the frequently updated trajectory remain close to the current state. In contrast, with delay, the interpolation strategy causes the setpoints to be farther from the current state, thereby enhancing the feedback effect.
We also investigated the effect of a second SQP iteration for $dt=25$ ms. Without delay, the number of successes increased by 13% at 30 ADMM iterations. However, delay makes the second SQP iteration ineffective, as success dropped to 0.02.